OpenStudy (anonymous):

7x^10=6x^5+5 in quadratic form

OpenStudy (mathstudent55):

Express the polynomial as a function of x^5. Do a substitution. Let u = x^5. Then 7x^10 = 7u^2. Now it's a quadratic in terms of u. Do the same with 6x^5.